OpenStudy (anonymous):

Consider the circle of radius 5 centered at (0,0). Find an equation of the line tangent to the circle at the point (3,4).

OpenStudy (anonymous):

would it help if i gave multiple choice?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

equation for the circle is \[x^2+y^2=5\] taking the derivative gives \[2x+2yy'=0\] \[2yy'=-2x\] \[y'=-\frac{x}{y}\] so slope at \[(3,4)\] is \[-\frac{3}{4}\]

OpenStudy (anonymous):

actually the equation for the circle is \[x^2+y^2=5^2\] \[x^2+y^2=25\] but that does not change the slope

sam (.sam.):

Hey satellite73, how you type the fraction sign?

OpenStudy (anonymous):

i dont understand. What is the answer?

OpenStudy (anonymous):